# Mind the Gap: Navigating Inference with Optimal Transport Maps

### v1

@@ Abstract
Machine learning (ML) techniques have recently enabled enormous gains in sensitivity across the sciences. In particle physics, much of this progress has relied on excellent simulations of a wide range of physical processes. However, due to the sophistication of modern machine learning (ML) algorithms and their reliance on high-quality training samples, discrepancies between simulation and experimental data can significantly limit the effectiveness of ML techniques. In this work, we present a solution to this “mis-specification” problem: a calibration approach based on optimal transport, which we apply to high-dimensional simulations for the first time. We demonstrate the performance of our approach through jet tagging, using a CMS-inspired dataset. A 128-dimensional internal jet representation from a powerful general-purpose classifier is studied; after calibrating this internal “latent” representation, we find that a wide variety of quantities derived from it for downstream tasks are also properly calibrated: using this calibrated high-dimensional representation, powerful new applications of jet flavor information can be utilized in LHC analyses. This is a key step toward allowing properly-calibrated “foundation models” in particle physics. More broadly, this calibration framework has broad applications for correcting high-dimensional simulations across the sciences.
@@ 1 Introduction
Model mis-specification occurs when the process of generating observations does not belong to the family of statistical models used for inference. This is a growing issue across disciplines: the physical sciences, economics, social sciences, medicine and health. As the use of inference techniques built on complex and precise statistical models, often enabled by sophisticated simulations, becomes more widespread, it is critical to develop techniques that mitigate mis-specification, enabling inference that is both effective and unbiased.
[…]
@@
Inference on approximately factorizable likelihoods is especially powerful. Scientific instruments are often designed to measure a quantity $x$ independent of the details of the particular experiment, aside from a latent variable $\omega$. For example, when measuring the weight of an object, an ideal scale is one that provides an accurate and precise estimate of the object’s weight, $p(x=\text{estimated weight}|\omega=\text{true weight})$, independently of whether the object is, say, a penguin or a banana ($\theta=\{\text{penguin},\text{banana}\}$). A scale with this factorization property can be calibrated independently using a set of test objects whose weights are well-known. In doing so, the experimenter constructs a good approximation of the scale’s density over outputs for a known input; this can be used to infer the weight of a variety of objects: penguins, bananas, or otherwise.
@@
In practice, likelihoods often do not factorize exclusively through some stochastic latent $\omega$ that can be measured in calibration experiments, but they depend on a variety of additional stochastic latents $\omega^{\prime}$. In such cases, it is not possible to derive the full dependence $p(x|\omega,\omega^{\prime})$, but it may still be possible to correct the explicit dependence on $\omega$.
@@
We are aware of two procedures for deriving these corrections to the simulated density, $p_{\text{sim}}(x|\omega)$, with calibration measurements, $p_{\text{data}}(x|\omega)$. The first constructs the density ratio
@@
[…]
Thus, corrections derived in the CR are applied in the SR via the weight $s_{\omega}(x)$; this is colloquially known as “vertical calibration”.
@@
[…]
@@
This “horizontal calibration” moves the simulated observations based on what is observed in the calibration region. The suitable transport map for correcting simulations is the optimal one, i.e. the one which minimally alters $p_{\text{sim}}(x|\omega)$ for each choice of $\omega$ while still closing with $p_{\text{data}}(x|\omega)$. The optimal transport (OT) map is solution to this closure problem that is least invasive, in the sense that it preserves the full structure of $p_{\text{sim}}(x|\omega,\omega^{\prime})$ as well as possible.
@@
It is worth noting that, in the limit that $\omega$ captures all the mismodeling of the simulation $p_{\text{sim}}(x|\omega,\omega^{\prime})$, then the the corrected simulation converges to the “true” likelihood $p_{\text{data}}$. This is true for both vertical and horizontal strategies. While it is not the focus of this article, the choice of $\omega$ against which one parameterizes the corrections is important; enabling corrections that are so conditioned on a maximally-informative $\omega$, of potentially very high-dimensions, is an active field of research.
@@
In the past several years, ML-based solutions to both the vertical and horizontal calibration problems have been developed. Density ratio estimation (DRE) via neural-network (NN) classifiers is a well-established technique that scales well to high dimensions in both $x$ and $\omega$ [CranmerEtAl2015]. More recently, neural OT solvers have been introduced [Original_ot_solution, cpflows, DBLP:journals/corr/abs-2106-01954, bunne2023supervisedtrainingconditionalmonge], enabling the derivation of optimal transport maps for horizontal calibrations of multi-dimensional simulations from which samples can be drawn [Pollard_2022, ATLAS-PAPER-2025].
@@
As we scale into higher dimensions, it is important to recognize that the horizontal and vertical strategies are distinct solutions to the calibration problem and exhibit different behaviors. In sample-based simulations, the reweighting function $s_{\omega}(x)$ used in vertical calibration will, in general, dilute the statistical power of the simulated samples by a factor
$\rho_{\omega}=\frac{1}{N}\frac{(\sum_{i}^{N}s_{\omega}(x_{i}))^{2}}{\sum_{i}^{N}s_{\omega}(x_{i})^{2}}.$ (5)
@@
Horizontal calibration does not introduce a dilution. Indeed, for well-behaved spaces (e.g. Euclidean space), a horizontal solution to a given calibration problem always exists; for vertical calibrations, a solution only exists if $\forall\,x\ \omega.\ p_{\text{data}}(x|\omega)>0\Rightarrow p_{\text{sim}}(x|\omega)>0$, i.e. the simulated distribution fully supports the data distribution and $s_{\omega}$ is always finite. These properties originate from the fact that an optimal transport distance between densities $p$ and $q$ over a space $Y$ is well-defined under rather loose conditions on $Y$ [villani2008optimal], while vertical calibrations inherently depend on the KL-divergence between $p$ and $q$ to be finite [Goodfellow-et-al-2016].
@@
In this article we explore a concrete example of an important calibration problem in particle physics. Data analysis in High Energy Physics (HEP) experiments at the Large Hadron Collider (LHC) aims to study the fundamental properties of matter by analyzing the complex final states resulting from high-energy collisions. These analyses rely on a hierarchical stochastic calculations, where simulations of the underlying “hard process,” involving fundamental particles, approximately factorize from the subsequent evolution into stable particles that eventually interact with the detector, whose signals undergo sequential processing through a series of algorithms defining the reconstruction chain of the experiment.
@@
Ultimately, a dedicated analysis is conducted to extract key parameters of the underlying theoretical models of interest, $\theta$, which govern the initial hard process. Within this framework, Monte Carlo (MC) simulations play a crucial role in modeling the likelihood via sampling. Although MC simulations are grounded in first-principles, they often exhibit substantial discrepancies when compared to experimental data, rendering $p_{\text{sim}}(x|\omega)$ unreliable at the required precision.
@@
Here we focus on an important use-case that is ubiquitous across the LHC physics programme: characterizing jets of hadronic particles (“jets”). For example, establishing what was the origin of a jet observed in a detector is critical for a broad number of important conclusions drawn from LHC data. ML techniques have vastly improved our ability to answer this question to high precision in recent years [UParT, GN2].
Recently, foundation models have emerged as promising tools for enhancing reconstruction tasks, summarizing the relevant features of jets in a general way [MPM, sophon, Birk_2024].
Foundation models used for jet tagging are essentially large-scale transformer-based networks that can be pre-trained to encode high-dimensional representations of jets, $z$. These latent representations may then be used as inputs for downstream inference tasks. While foundation models offer significant potential for enhancing HEP reconstruction pipelines, experimental methodologies for calibrating discrepancies between MC simulations and real experimental data in the latent representation are currently lacking. In this work, we build upon previous research and, for the first time, introduce a calibration approach leveraging optimal transport theory, applied directly in a high-dimensional latent space of up to 128 dimensions. Calibration algorithms based on optimal transport have been introduced in Ref. [Pollard_2022, ATLAS-PAPER-2025] and applied to correct low-dimensional data representations, such as the output scores of classifiers used for jet tagging in Ref. [GN2, ATLAS:2019EPJC79-970]. While we demonstrate this methodology in the context of jet tagging, its applicability extends to a wide range of reconstruction tasks in HEP experiments.
This article describes two primary contributions in this direction. First, we find that density ratio-based calibrations are not suitable in our real-world case, because the densities between our base model and simulated data are not sufficiently overlapping. Second, we find that optimal-transport solutions to the calibration problem scale up to high dimensionality and properly close in an inference-relevant way.
@@ 2 Related work
Optimal transport (OT) has recently emerged as a powerful tool for constructing high dimensional maps between distributions, with a growing body of applications in both theoretical and experimental physics [brenier2002cosmo, levy2022monge, vonhausegger2021bao, otsham2024].
In high-energy physics (HEP), one of the primary applications of OT is to correct discrepancies in simulated datasets through calibration [Pollard_2022]. This approach has been adopted by the ATLAS experiment, which employed OT to calibrate algorithms used for $b$-jet identification [ATLAS-PAPER-2025]. To the best of the author’s knowledge, this represents the first instance of a horizontal calibration method applied to jet tagging. In contrast, conventional jet-tagging calibration techniques predominantly rely on vertical calibration strategies [ATLAS:2019EPJC79-970, CMS:2020P06005, ATLAS-PHYS-PUB-2021-035, ATLAS-CONF-2017-064, CMS:2022P03014].
OT has also been explored in the modeling of complex background processes in the context of double Higgs bosons production at the LHC [manole2022background], for anomaly detection algorithms [craig2024anomaly], and for variable decorrelation [Algren_2024].
These advances are grounded in foundational developments by Monge and Kantorovich [villani2008optimal], as well as in algorithmic contributions that enabled scalable applications of OT in machine learning [peyre2019computational].
@@ 3 Simulated dataset
The dataset utilized in this study is based on JetClass [JetClass], a well-established benchmark for jet classification tasks in HEP. It is generated using standard Monte Carlo event generators widely adopted in LHC experiments and incorporates detector response and reconstruction effects simulated with DELPHES [delphes] (v3.4.2).
Jet reconstruction is performed using the anti-$k_{T}$ clustering algorithm [Cacciari_2008, Cacciari_2012] with a radius parameter of $R=0.8$, applied to the E-Flow objects reconstructed by DELPHES. The analysis focuses on jets with transverse momenta in the range of 5005005005001000100010001000 GeV and pseudorapidity satisfying $|\eta|<2$.
Each jet in the dataset is assigned to one of ten distinct classes ($l$), corresponding to the underlying physics processes responsible for its production. These classes include Higgs boson decays ($H\rightarrow b\bar{b}$, $H\rightarrow c\bar{c}$, $H\rightarrow gg$, $H\rightarrow 4q$), top quark decays ($t\rightarrow bqq^{\prime}$, $t\rightarrow b\ell\nu$), hadronic decays of $W$ and $Z$ bosons, and a background class labeled $q/g$, which represents jets originating from light quarks ($u$, $d$, $s$) or gluons ($g$): $l\in{Hbb,Hcc,Hgg,H4q,Hl\nu qq,tbqq^{\prime},tbl\nu,Wqq,Zqq,q/g}$.
This study focuses on the calibration of $q/g$ jets which are abundantly produced at the LHC. The samples are generated using QCD multijet processes simulated with Pythia v8.186 [Sj_strand_2015]. Parton showering and hadronization are also modeled with the same generator. Additionally, multiple parton interactions are included in the simulation, while pile-up effects are not considered.
To emulate the typical discrepancies observed between simulated samples and real collision data in LHC experiments, two datasets of QCD multijet events are generated. The first dataset, referred to as the source, is produced using the standard JetClass configuration. The second dataset, referred to as the target, is created by modifying the smearing functions of the DELPHES detector simulation. In particular, the impact parameters of charged particles are additionally smeared by 10%–30%, based on informed estimates from Ref. [ATLAS-CONF-2018-006]. In parallel, the renormalization and factorization scales of the QCD multijet events are increased from 1 to 2. These modifications are intended to simulate two main sources of systematic differences: inaccuracies in the modeling of the underlying physical processes, and effects stemming from detector response and reconstruction. Figure 1 compares the distributions of the transverse impact parameter ($d_{0}$) and the number of jet constituents for the two datasets.
[…]
ıin csts_input_d0val, jets_jet_nparticles

### v2

@@ Abstract
Machine learning (ML) techniques have recently enabled enormous gains in sensitivity to new phenomena across the sciences. In particle physics, much of this progress has relied on excellent simulations of a wide range of physical processes. However, due to the sophistication of modern machine learning algorithms and their reliance on high-quality training samples, discrepancies between simulation and experimental data can significantly limit their effectiveness. In this work, we present a solution to this “misspecification” problem: a model calibration approach based on optimal transport, which we apply to high-dimensional simulations for the first time. We demonstrate the performance of our approach through jet tagging, using a dataset inspired by the CMS experiment at the Large Hadron Collider. A 128-dimensional internal jet representation from a powerful general-purpose classifier is studied; after calibrating this internal “latent” representation, we find that a wide variety of quantities derived from it for downstream tasks are also properly calibrated: using this calibrated high-dimensional representation, powerful new applications of jet flavor information can be utilized in LHC analyses. This is a key step toward allowing the unbiased use of “foundation models” in particle physics. More broadly, this calibration framework has broad applications for correcting high-dimensional simulations across the sciences.
@@ 1 Introduction
The use of inference techniques built on complex and precise statistical models, often enabled by sophisticated simulations, continues to become more widespread across the sciences. This has led to a recent boom of so-called “simulation-based inference“ (SBI), in which neural methods are used to perform inference over information-rich observations [1, 2, 3, 4]. When the family of statistical models used for inference does not cover the generation process of observed data, a misspecification gap emerges, and constraints on model parameters will no longer be robust. Indeed, there is a growing body of work quantifying the impact of misspecification in a variety of environments [5] and in particular in the context of SBI [6, 7, 8]. Given the power of neural methods to gracefully handle high-dimensional and informative observations, it is critical to develop techniques that mitigate misspecification, enabling inference that is both effective and unbiased.
In the presence of control data, optimal transport (OT) theory [9, 10] provides a framework for addressing the misspecification problem through model calibration. Pollard and Windischhofer [11] argued that OT is an appropriate solution to model misspecification, and the ATLAS collaboration [12] successfully applied this strategy to calibrate three-dimensional simulations with particle collider data from the Large Hadron Collider [13]. A similar approach was subsequently applied to calibrate models for cardiovascular health [14].
In this work, we show both in simplistic and realistic examples that the OT-based calibration procedure scales well to high dimensions for an important class of hierarchical models. We find that density ratio-based corrections to samples drawn from a forward model are unsuitable, even in simple cases, for a large body of high-dimensional problems. We additionally show that, in a familiar scenario from particle physics, the densities between a baseline forward model and simulated data are not sufficiently overlapping for density-ratio corrections to be feasible. Moreover, we derive OT solutions to address the misspecification gap in the same particle-physics example and conclude that the model misspecification is adequately corrected in an inference-relevant way. This result lays the groundwork for properly-calibrated foundation models in scenarios with suitable calibration data, which enables their use for high-performance, but unbiased, statistical inference.
2 Calibration of hierarchical models
In hierarchical models, the likelihood of observing some datum $x$, given a parameter of interest $\theta$, is determined by marginalizing over any stochastic latent variable $\omega$: $p(x|\theta)=\int d\omega\,p(x|\omega,\theta)\,p(\omega|\theta)$. An important subset of hierarchical models comprises those which factorize, such that $x$ only depends on $\theta$ through $\omega$, i.e. $p(x|\omega,\theta)=p(x|\omega)$. In such situations, it may be possible to calculate $p(x|\omega)$ separately, or to constrain it via auxiliary experiments. Indeed, this is often the motivation for constructing the likelihood $p(x|\theta)$ in terms of a stochastic latent $\omega$: when $p(x|\theta)\approx\int d\omega\,p(x|\omega)\,p(\omega|\theta)$, then both $p(x|\omega)$ and $p(\omega|\theta)$ can be independently derived or verified.
@@ 2 Calibration of hierarchical models
Inference on approximately factorizable likelihoods is especially powerful. Scientific instruments are often designed to measure a quantity $x$ regardless of the details of the particular experiment, except for the stochastic latent variable $\omega$. For example, when measuring the mass of an object, an ideal scale is one that provides an accurate and precise estimate of the object’s weight, $p(x=\text{estimated mass}\,|\,\omega=\text{true mass})$, independently of whether the object is, say, a penguin or a banana. A scale with this factorization property can be calibrated using a set of test objects whose weights are well-known; in doing so, the experimenter constructs a good approximation of the scale’s density over outputs for a known input. An object’s mass can then be used to faithfully infer the identity of the object, ($\theta\in\{\text{penguin},\text{banana}\}$).
@@
In practice, likelihoods often do not factorize exclusively through some $\omega$ that can be measured in calibration experiments, but they depend on a variety of additional stochastic latents $\omega^{\prime}$. In such cases, it is often not possible to derive the full dependence $p(x|\omega,\omega^{\prime})$, but it may still be possible to correct the explicit dependence on $\omega$ and remove the likelihood “gap” between reality and the forward model used for inference.
2.1 Calibration methods
@@
We are aware of two procedures for deriving these corrections to the simulated density, $p_{\text{sim}}(x|\omega)$, with calibration data drawn from $p_{\text{data}}(x|\omega)$. The first constructs the density ratio
@@
$p(x|\theta)=\int d\omega\,d\omega^{\prime}\,s_{\omega}(x)\,p_{\text{sim}}(x|\omega,\omega^{\prime})\,p_{\text{sim}}(\omega,\omega^{\prime}|\theta).$ (2)
Thus, corrections derived in the CR are applied to the SR likelihood via the weight $s_{\omega}(x)$; this is colloquially known as “vertical calibration”.
@@
$p(x|\theta)=\int d\omega\,d\omega^{\prime}\,(T_{\omega})_{\#}\,p_{\text{sim}}(x|\omega,\omega^{\prime})\,p_{\text{sim}}(\omega,\omega^{\prime}|\theta).$ (4)
@@
This “horizontal calibration” moves the simulated observations based on what is observed in the calibration region. The suitable transport map for correcting simulations is the optimal one, i.e. the one which minimally alters $p_{\text{sim}}(x|\omega)$ for each choice of $\omega$ while still closing with $p_{\text{data}}(x|\omega)$ [11]. The optimal transport map is the solution to this closure problem that is least invasive, in the sense that it preserves the full structure of $p_{\text{sim}}(x|\omega,\omega^{\prime})$ as well as possible.
@@
It is worth noting that, in the limit that $\omega$ captures all the mismodeling of the simulation $p_{\text{sim}}(x|\omega,\omega^{\prime})$, then the the corrected simulation converges to the “real” likelihood $p_{\text{data}}$. This is true for both vertical and horizontal strategies. While it is not the focus of this article, the choice of $\omega$ against which one parameterizes the corrections is important; enabling corrections that are conditioned on a sufficiently informative $\omega$, of potentially very high-dimensions, is an active field of research.
@@
In the past several years, ML-based solutions to both the vertical and horizontal calibration problems have been developed. Density ratio estimation (DRE) via neural-network (NN) classifiers is a well-established technique that scales well to high dimensions in both $x$ and $\omega$. More recently, neural OT solvers have been introduced [15, 16, 17, 18], enabling the derivation of OT maps for horizontal calibrations of multi-dimensional simulations from which samples can be drawn [11, 12].
@@
As we scale into higher dimensions, it is important to recognize that the horizontal and vertical strategies are distinct solutions to the misspecification problem and exhibit different behaviors. In sample-based simulations, the reweighting function $s(x)$ used in vertical calibration will, in general, dilute the statistical power of the simulated samples by a factor
$\rho\equiv\frac{\mathbb{E}_{x\sim p_{\text{sim}}}[s(x)^{2}]}{\mathbb{E}_{x\sim p_{\text{sim}}}[s(x)]^{2}}=\mathbb{E}_{x\sim p_{\text{data}}}[s(x)],$ (5)
i.e. the effective number of samples available after calibration is $N_{\mathrm{eff}}\equiv N_{\mathrm{sim}}/\rho$, where $N_{\mathrm{sim}}$ is the original number of samples. Appendix A shows the derivation of this equality in more detail. Horizontal calibration does not introduce a dilution.
@@
Indeed, for well-behaved spaces (e.g. Euclidean space), a horizontal solution to a given calibration problem always exists. For vertical calibrations, a solution only exists under certain conditions; at the very least, $\forall\,x\ \omega.\ p_{\text{data}}(x|\omega)>0\Rightarrow p_{\text{sim}}(x|\omega)>0$ must hold, i.e. the simulated distribution must fully support the data distribution such that $s_{\omega}$ is always finite. However, as shown in Appendix A, this is not in general sufficient to guarantee existence of a vertical-calibration solution [^1] For a more in-depth discussion of importance sampling weights and their behaviors in a variety of situations, including potential pitfalls, we recommend Chapter 9 of Owen’s Monte Carlo: Theory, Methods and Examples [19] or Chapter 3 from Robert & Casella [20].. These properties originate from the fact that an OT distance between densities $p$ and $q$ over a space $Y$ is well-defined under rather loose conditions on $Y$ [10], while vertical calibrations inherently depend on the $\chi^{2}$-divergence between $p$ and $q$ to be finite [21, 22].
As derived in Appendix A, even in simple scenarios in which $p_{\text{data}}$ and $p_{\text{sim}}$ are approximately multivariate-normal distributed (MVN), there are stringent requirements on how overlapping they must be for vertical calibrations to be viable. In the MVN case, if the variance of $p_{\text{data}}$ is at least twice that of $p_{\text{sim}}$ in even one dimension, then the dilution factor is infinite, and vertical calibrations are not possible. Even if the covariances are identical, the dilution factor grows exponentially with the Mahalanobis distance [23] of the means of $p_{\text{data}}$ and $p_{\text{sim}}$, again limiting the utility of the vertical strategy. Indeed, in the particle physics-inspired example explored in Section 6, we conclude that $\rho$ is very large, and a vertical calibration solution may not exist.
2.2 Hierarchical models in collider physics
@@
In this article we explore a concrete example of an important calibration problem in particle physics. Data analysis in High Energy Physics (HEP) experiments at the Large Hadron Collider (LHC) aims to study the fundamental properties of matter by analyzing the complex final states resulting from high-energy collisions [13]. These analyses rely on hierarchical stochastic calculations, where simulations of the underlying “hard process,” involving fundamental particles, approximately factorize from the subsequent evolution into stable particles that eventually interact with the detector. These detector signals undergo sequential processing through a series of algorithms defining the reconstruction chain of the experiment [24, 25, 26, 27, 28].
@@
Ultimately, dedicated analyses are conducted to extract key parameters of the underlying theoretical models of interest, $\theta$, which govern the initial hard process. Within this framework, Monte Carlo (MC) simulations play a crucial role in modeling the likelihood via sampling. Although MC simulations are grounded in first-principles, they often exhibit substantial discrepancies when compared to experimental data, rendering $p_{\text{sim}}(x|\omega,\theta)$ unreliable at the required precision.
@@
2.3 Hadronic jets and their representations
Here we focus on an important use-case that is ubiquitous across the LHC physics programme: characterizing jets of hadronic particles (“jets”). For example, establishing what was the origin of a jet observed in a detector is critical for a broad number of important conclusions drawn from LHC data. ML techniques have vastly improved our ability to answer this question to high precision in the last few years [29, 30].
Even more recently, foundation models have emerged as promising tools for enhancing reconstruction tasks, summarizing the relevant features of jets in a general way [31, 32, 33, 34, 35]. Foundation models used for jet tagging are essentially large-scale transformer-based networks that can be pre-trained to encode high-dimensional representations of jets, $z$. These latent representations may then be used as inputs for downstream inference tasks. While foundation models offer significant potential for enhancing HEP reconstruction pipelines, experimental methodologies for calibrating discrepancies between MC simulations and real experimental data in the latent representation are currently lacking. In this work, we build upon previous research and, for the first time, introduce a calibration approach leveraging OT theory, applied directly in a high-dimensional latent space of up to 128 dimensions. Calibration algorithms based on optimal transport have been introduced in Ref. [11, 12] and applied to correct low-dimensional data representations, such as the output scores of classifiers used for jet tagging in Ref. [30, 36]. We demonstrate this methodology’s promise in addressing misspecification for a much more informative set of features. While we concentrate on the objective of jet tagging, the applicability of this technique extends to a wide range of reconstruction tasks in HEP experiments and beyond.
@@ 3 Related work
Optimal transport has recently emerged as a powerful tool for constructing high dimensional maps between distributions, with a growing body of applications in both theoretical and experimental physics [37, 38, 39, 40].
In high-energy physics (HEP), one of the primary applications of OT is to correct discrepancies in simulated datasets through calibration [11]. This approach has been adopted by the ATLAS experiment, which employed OT to calibrate algorithms used for $b$-jet identification [12]. To the best of the author’s knowledge, this represents the first instance of a horizontal calibration method applied to jet tagging. In contrast, conventional jet-tagging calibration techniques predominantly rely on vertical calibration strategies [36, 41, 42, 43, 44]. Alternative horizontal calibration methods have been proposed, e.g. via normalizing flows [45], but it is important to recognize that these solutions are known not to converge on the optimal-transport map in real-world situations [46], nor are the solutions unique.
OT has also been explored in the modeling of complex background processes in the context of double Higgs bosons production at the LHC [47], for anomaly detection algorithms [48], and for variable decorrelation [46].
These advances are grounded in foundational developments by Monge and Kantorovich [10], as well as in algorithmic contributions that enabled scalable applications of OT in machine learning [49].
@@ 4 Simulated dataset
The dataset utilized in this study is based on JetClass [50], a well-established benchmark for jet classification tasks in HEP. It is generated using standard Monte Carlo event generators widely adopted in LHC experiments and incorporates detector response and reconstruction effects simulated with DELPHES [51] (v3.4.2).
Jet reconstruction is performed using the anti-$k_{T}$ clustering algorithm [52, 53] with a radius parameter of $R=0.8$, applied to the E-Flow objects reconstructed by DELPHES. The analysis focuses on jets with transverse momenta in the range of 50050010001000 GeV and pseudorapidity satisfying $|\eta|<2$.
Each jet in the dataset is categorized into one of ten distinct classes ($l$), each reflecting the underlying physics process responsible for its production. The classes comprise Higgs boson decays ($H\rightarrow b\bar{b}$, $H\rightarrow c\bar{c}$, $H\rightarrow gg$, $H\rightarrow 4q$), top quark decays ($t\rightarrow bqq^{\prime}$, $t\rightarrow b\ell\nu$), hadronic decays of the $W$ and $Z$ bosons, and a background class denoted $q/g$, which corresponds to jets initiated by quarks or gluons not included in the previous categories:
$l\in{Hbb,Hcc,Hgg,H4q,Hl\nu qq,tbqq^{\prime},tbl\nu,Wqq,Zqq,q/g}$.
This study focuses on the calibration of $q/g$ jets which are abundantly produced at the LHC. The samples are generated using QCD multijet processes simulated with Pythia v8.186 [54]. Parton showering and hadronization are also modeled with the same generator. Additionally, multiple parton interactions are included in the simulation, while pile-up effects are not considered.
To emulate the typical discrepancies observed between simulated samples and real collision data in LHC experiments, two datasets of QCD multijet events are generated. The first dataset, referred to as the source, is produced using the standard JetClass configuration. The second dataset, referred to as the target, is created by modifying the smearing functions of the DELPHES detector simulation. In particular, the impact parameters of charged particles are additionally smeared by 10%–30%, based on informed estimates from Ref. [55]. In parallel, the renormalization and factorization scales of the QCD multijet events are increased from 1 to 2. These modifications are intended to simulate two main sources of systematic differences: inaccuracies in the modeling of the underlying physical processes, and effects stemming from detector response and reconstruction. Figure 1 compares the distributions of the transverse impact parameter ($d_{0}$) and the number of jet constituents for the two datasets.